How in the world do you solve this rational equation using cross-multiply/ Lowest Common Denominator (LCD):
1/2x + 1/4x = 7/8x
1 over 2x PLUS 1 over 4x EQUALS 7 over 8x

Answer by Alan3354(69443)

1/2x + 1/4x = 7/8x
2/4x + 1/4x = 7/8x
3/4x = 7/8x
3*8x = 7*4x
24x = 28x
-4x = 0
x = 0
----------
But that makes all the denominators zero, so
No solution.